OpenStudy (anonymous):

A certain plane can fly 210 mph in calm air. if the plane can travel 540 miles against the wind in the same amount of time it can travel 810 miles with the wind, then what is the rate of wind?

OpenStudy (anonymous):

with wind speed = 210 + x against speed = 210-x time = distance / speed t = 810 / 210 + t t= 540 / 210 - r 810 / 210+x = 540 / 210 - x 810(210 - x) = 540(210 + x) solve this to get wind speed

OpenStudy (anonymous):

that comes to 42 mph

OpenStudy (rulnick):

210 t = 675 (average of 540 and 810), to t=675/210 (= about 3.214) hours We'll use t=675/210 below. (210+x) t = 810 (210-x) t = 540 --------------------- 2xt = 270 x = 135/t So x=135/(675/210) = 42 mph.

OpenStudy (rulnick):

So a plane of speed 210 mph can travel 540 miles against the wind or 810 miles with the wind, when wind speed is 42 mph.